Circular No. 6639

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997by IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6418) reports the discovery of an apparent supernova (mag V = 16.2) on V and R CCD images taken on Apr. 27 by S. Chan on the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1997by is located near Abell 1736 at α = 13^{h}27^{m}27^{s}.54, δ = −27^{o}19′21′′.7 (equinox 2000.0). A nearby star has position end figures 18^{s}.62, 18′54′′.0. SN 1997by was not detected on frames taken on Apr. 4, but it is clearly visible on CCD images taken with the Mount Stromlo 0.76-m telescope on Apr. 28. Finding charts are available

at http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

2EG J0432+2910, 2EG J0744+5438, 2EG J0852-1237 J. P. Halpern and M. Eracleous, University of California at Berkeley, re-

port optical identifications based on spectrophotometry with the Kitt Peak 2.1-m telescope of flat-spectrum radio sources that have been suggested as EGRET blazar counterparts: “The radio and optical object in the error box of 2EG J0432+2910 (located at α = 4^{h}33^{m}37^{s}.8, δ = +29^{o}05′56′′, equinox 2000.0), described by Lundgren et al. (IAUC 6258), and Condon et al. (1983, A.J. 88, 20) has a very red BL Lac-type continuum spectrum with slope −4.0 and no emission or absorption features in the range 390–750 nm to an upper limit of 0.2 nm equivalent width (EW); from the spectrum obtained on 1997 Feb. 6, it has V = 17.8, which is ∼ 2 mag brighter than on the Palomar Sky Survey. The radio source candidate for 2EG J0744+5438 (located at α = 7^{h}42^{m}39^{s}.8, δ = +55^{o}44′25′′) proposed by Mukherjee et al. (1995, Ap.J. 445, 189) is also highly variable in the optical, having declined from V = 16.9 on 1996 Feb. 16 to V = 18.4 on 1997 Feb. 6; it has z = 0.723, based on weak emission lines in the low-state spectrum that we identify with Mg II 280 nm (EW = 2.0 nm), Hγ, and Hβ, and its slope is −1.0 in the low state and −1.3 in the high state. If this identification is correct, it would be the weakest radio counterpart of an EGRET blazar, with a 5-GHz flux density of 0.27 Jy. The candidate for 2EG J0852–1237 (located at α = 8^{h}50^{m}09^{s}.6, δ = −12^{o}13′34′′) described by Mattox et al. (1997, Ap.J., May 20 issue) is also highly variable, having brightened from V = 18.9 on 1996 Feb. 16 to V = 16.8 on 1997 Feb. 6. Its redshift is z = 0.566 from the same emission lines as above, plus [O III] 500.7-nm. In the high-state spectrum, Mg II 280-nm has EW = 1.0 nm and the continuum slope is −1.2.”

1997 April 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green